what would be the total investment and the interest earned when $2,500 is invested at 5% compounded semiannually for 3 years?
Answer by stanbon(75887)

what would be the total investment and the interest earned when $2,500 is invested at 5% compounded semiannually for 3 years?
----
A(t) = P(1+(r/n))^(nt)
A(3) = 2500(1+(0.05/2))^(2*3)
---
A(3) = $2899.23
===================
Interest = 2899.23-2500 = $399.23
-------
Cheers,
Stan H.